OpenStudy (rayman5766):

Variable y varies directly with variable x, and y = 5 when x = 8. What is the value of x when y = 15?

OpenStudy (rayman5766):

@iiForeverLost alright smart guys lets see if you can do that again

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Like for example the last problem was like this: If they vary directly then y = kx. Plug in the two values to find k. y = kx 6 = k(20) k = 6/20 = 3/10 = 0.3

OpenStudy (anonymous):

If im Correct So What is your Y Do you know?

OpenStudy (anonymous):

Ok.. It Would be kx Like last Time So Lets Pretend 6 Is x ok? Now What does k = and Then You Divide that fraction By 2 Then Covert To a decimal

OpenStudy (rayman5766):

yea don't think I am stupid but my teacher gave me this and I was late for class so I missed most of what is going so if I don I hope you understand

OpenStudy (anonymous):

Mhm Well Basicly you do that top your Probolem again just diffrnent numbers Ok?

OpenStudy (rayman5766):

pluse I am not good with word problems be like uhh

OpenStudy (rayman5766):

also I did not mean to e at plus also why did you give me a medal I dint solve it yet

OpenStudy (anonymous):

Ok.. The answer should be about 8.33333333333(Ect)

OpenStudy (rayman5766):

so bascule 8.3

OpenStudy (anonymous):

Yea